OpenStudy (anonymous):

f(x)=0. show that there is a root for equation: sin(x)=x^2-x for 1

OpenStudy (anonymous):

We know \[-1\leq \sin x \leq 1\] hence, \[-1\leq -\sin x \leq 1\] Now since, \[1<x<2\] and hence, \[-2<-x<-1\] Therefore\[1<x^2<4\] We can rewrite the equation as \[x^2-x-\sin x = 0\] Therefore adding equations above we get \[-2<x^2-x-\sin x < 4\] Therefore for some value of x the equation has a very equals to 0 and hence a root exists.

OpenStudy (anonymous):

Is this a question that you have in calculus?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let f(x) = sinx - x^2 + x a special case of the intermediate value theorem states: if 0 is a number between f(1) and f(2), then there is a c in [1, 2] such that f(c) = 0 (ie: f has a root) f(1) = sin(1) - 1 + 1 = sin(1), which is greater than 0 f(2) = sin(2) - 4 + 2 = sin(2) - 2, which is less than 0 hence there is a c in [1,2] such that f(c) = 0

OpenStudy (anonymous):

technically the inverval should be (1,2) , ie not including the endpoints :P

OpenStudy (anonymous):

but if there is a c in [1,2] such that f(c) = 0, that implies there is a c in (1,2) since f(1) and f(2) aren't 0